# Leveraging Large Language Models to Address Data Scarcity in Machine Learning: Applications in Graphene Synthesis

### v1

@@ Abstract
Machine learning in materials science faces challenges due to limited experimental data, as generating synthesis data is costly and time-consuming, especially with in-house experiments. Mining data from existing literature offers a potential solution but introduces issues like mixed data quality, inconsistent formats, and variations in reporting experimental parameters, complicating the creation of consistent features for the learning algorithm. Additionally, combining continuous and discrete features can hinder the learning process with limited data. Here, we propose strategies that utilize large language models (LLMs) to enhance machine learning performance on a limited, heterogeneous dataset of graphene chemical vapor deposition synthesis compiled from existing literature. These strategies include prompting modalities for imputing missing data points and leveraging large language model embeddings to encode the complex nomenclature of substrates reported in chemical vapor deposition experiments. Imputed data from various prompting modalities are compared to statistical K-nearest neighbors imputation, with the former yielding a more diverse distribution, richer feature representation, and improved model generalization. We also demonstrate that discretizing continuous features enhances prediction accuracy. The proposed strategies enhance graphene layer classification using a support vector machine (SVM) model, increasing binary classification accuracy from 39% to 65% and ternary accuracy from 52% to 72%. We compare the performance of the SVM and a GPT-4 model, both trained and fine-tuned on the same data. Our results demonstrate that the numerical classifier, when combined with LLM-driven data enhancements, outperforms the standalone LLM predictor, highlighting that in data-scarce scenarios, improving predictive learning with LLM strategies requires more than simple fine-tuning on datasets. Instead, it necessitates sophisticated approaches for data imputation and feature space homogenization to achieve optimal performance. The proposed strategies emphasize data enhancement techniques rather than merely refining learning architectures or regularizing loss functions, offering a broadly applicable framework for improving machine learning performance on scarce, inhomogeneous datasets.
@@ 1 Introduction
Large language models (LLMs) have recently emerged as a new approach in predictive chemistry and materials science 37, 38, 39, 40, 41, 42, 43, 44, 45. Specifically, they hold significant potential to address data scarcity and heterogeneity challenges, particularly by utilizing the extensive embedded pre-trained knowledge to impute missing data or homogenize inconsistencies in reported features. LLMs are referred to as near-universal generalists 46, positioning them as readily accessible alternatives to the aforementioned strategies in the field of materials informatics. Furthermore, LLMs can be fine-tuned for specific problems with minimal prompt engineering, making them technically feasible for navigating complex feature spaces 47. This flexibility and generality offer potential solutions to both underfitting and overfitting challenges in machine learning on small datasets, while also enabling LLMs to minimize hallucinations during data generation or featurization. In essence, LLMs could help address the bias-variance tradeoff that most ML strategies struggle to resolve when working with limited data. In this study, we demonstrate an LLM-assisted enhancement of machine learning using a sparsely populated and inherently heterogeneous dataset on chemical vapor deposition (CVD) of graphene, compiled from multiple studies in the literature.
@@
The CVD synthesis of graphene involves a wide range of experimental setups, with several key parameters governing the growth of graphene layers 48, 49, 50. To date, no comprehensive dataset encompassing an exhaustive set of synthesis parameters has been developed for graphene CVD growth or for synthesizing any 2D (atomic-layer-thick) material. Consequently, most machine learning studies on 2D material growth rely on limited, in-house experimental data, often constrained to specific ranges of synthesis conditions (e.g., pressures, temperatures, and precursor flow rates). These studies typically utilize machine learning models that require minimal feature engineering, as the in-house experimental data is already labeled and well-structured 51, 52, 53, 54, 55. For example, Hwang et al. developed a machine learning model to predict the morphology of graphene in the form of SEM (Scanning Electron Microscopy) images based on CVD processing variables (i.e., temperature, annealing time, growth time, and supply of hydrogen) 55. The size, coverage, domain density, and aspect ratio of CVD-grown graphene in SEM images were automatically measured using a region proposal convolutional neural network (R-CNN). A neural network and support vector machine were trained to map the relationship between CVD processing variables and these morphological features. Subsequently, a generative adversarial network (GAN) was designed to generate SEM images based on specified CVD processing conditions 55. However, this study relied on in-house experimental data, restricted to four processing variables and limited by fixed chamber conditions, including a constant precursor flow rate and the exclusive use of a single substrate type (Cu). Another notable effort by Schiller et al. involved creating a crowd-sourced database compiling an extensive list of parameters influencing various aspects of graphene CVD growth, contributed by multiple experimental groups 15. They also developed a UNet model for SEM image characterization, accessible as a nanoHub tool 56. However, the database remains incomplete, making it challenging to integrate into a machine learning framework for predicting outcomes based on CVD processing conditions.
@@
Recent advancements in large language models enable them to perform complex tasks such as understanding context and semantics and generating contextually relevant responses. The vast training data, pretraining on diverse corpora, and the sheer scale of these models make them excel at interpreting meaning and delivering accurate, coherent outputs. Several studies have utilized LLMs for imputation tasks 65, 66, 67, 68, 69, 70, 71, showing superior performance compared to traditional techniques such as K nearest neighbor (KNN), mean imputation, multivariate imputation by chained equations (MICE) and ML methods such as generative and discriminative methods (GAN) 72, 70, 66, 73. For example, Nazir et al. utilized ChatGPT-3.5 as an imputation engine to populate missing values for biological and psychological data, showcasing its out-performance compared to KNN and mean imputation65. They applied iterative human-in-the-loop feedback process to refine ChatGPT’s imputation response by comparing the preceding response with ground-truth data, resulting in context-aware and accurate imputations of missing values in subsequent steps. Another study by Hayat et al.66 introduced an LLM-based imputation formalism to populate missing values for seven multi-variate classification datasets from the UCI repository74. They generated missing-value descriptors using a pre-trained LLaMA2 model and then fine-tuned the model using the descriptor-enriched data set for classification tasks. The classification tasks were directly performed using the fine-tuned LLMs. Other notable studies have incorporated techniques such as QnA prompts for data imputation and property extraction using LLMs across a wide variety of tasks 73, 75, 72. This study focuses on leveraging the existing pre-trained ChatGPT-4o-mini for a general imputation strategy. Additionally, the learning capacities of numerical models, such as support vector machines and random forests, are evaluated for the target classification task, and are compared with a fine-tuned ChatGPT-4o-mini counterpart.
@@
LLM-based data imputation is benchmarked against KNN-based imputation for the classification task. For the LLM-based approach, various prompt-engineering strategies are employed to guide the imputation task with specific instructions. After imputation, LLM-based featurization is used to unify and homogenize the numerical features for training binary and ternary classifiers, which predict the number of graphene layers based on a set of CVD processing inputs. We compare the effectiveness of different prompt-engineering strategies with traditional KNN-based data imputation, highlighting the superior performance of specific prompt-engineered GPT models. The prompt-driven imputed values are benchmarked with existing ground truth, and the missing values are directly imputed as an independent task. LLM-based strategies minimize imputation errors while enhancing data diversity, yielding a knowledge-driven, less imitative data distribution that outperforms KNN in machine learning. Additionally, we compare traditional non-LLM label-encoded features, commonly used for handling inconsistent labels, with LLM-generated features. Our results demonstrate the superior performance of LLM-embedded features in enhancing ML performance compared to label- or category-encoded features. Finally, we discretize the attributes to enhance the model’s learning performance compared to their continuous counterparts. These techniques collectively improve ML performance for graphene layer classification, as detailed in the following sections.
@@ 2.1 Data Collection
Experimental literature was used to manually extract 164 data entries 57, 58, 59, 60, 50, 61, 62, 63, 64, each consisting of 10 attributes. Five of these attributes represent precursor flow rates (sccm), while the remaining chamber parameters include pressure (mbar), temperature (${}^{\circ}\text{C}$), growth time (min), and substrate. The target variable is the number of graphene layers. Many studies also report the CVD method; however, this primarily reflects the range of pressure used in the CVD process. Therefore, we did not include it as a separate attribute. Table 1 shows a subset of the dataset used in this study. The dataset contains missing values for pressure, growth time, number of layers, and nearly all precursor flow rates. The precursor flow rates for each data entry are presented in the graphene_growth_conditions_layers.csv supplementary file. The distributions of raw data across various attributes are shown in Supplementary Figure 1. More details on data collection are provided in Supplementary Note 1.
@@ 2.2 Imputation of Missing Attributes
Besides KNN imputation, we employed prompt-based LLM imputations to better handle sparsely populated or nearly empty attributes, such as the gas flow rates, which did not provide a well-defined data distribution for a distance-based algorithm like KNN. A more “generally-informed” model, such as an LLM, can leverage syntactical relationships between words to generate data that closely approximates values that might have been reported in the literature. Figure 2 shows the prompt-based LLM imputation methodology developed in this study. The prompts used for this task are named based on their general purpose as summarized in Table 2. Supplementary Note 1.1 details the methodology for each prompt. The GIST prompt introduced here is inspired by and an approximation of the CLAIM prompt approach introduced in Ref.66. GIST differs from CLAIM as it only incorporates the language-descriptors as a pre-processing step to assist the imputation and does not create a missingness-aware description of the data which is central to the CLAIM methodology. Identifying missingness patterns is not feasible with a high-dimensional, low-cardinality data like ours. The imputed datasets with each method are carried onto the next step of feature engineering. Supplementary Notes 1.3 to 1.5 provide details on imputation methods and their benchmarking.
@@
where LLM model = ChatGPT-4o-mini, context vector = [attributes] - [target], attribute = current attribute being imputed. We also demonstrate a benchmarking experiment where ChatGPT imitates the KNN imputation approach, shown in Supplementary Figure 3. Additional details about the nature of the imputed data points is shown in Supplementary Figures 4 and 5.
@@ 2.3 Feature Engineering
We conduct feature engineering to process the substrate attribute and the target variable, the number of layers. The latter is handled using label encoding, classifying graphene layers into binary categories (monolayer versus multilayer) or ternary categories (monolayer versus bilayer versus multilayer), as the values are discrete. However, for the substrate attribute, as shown in Table 1, the data was inconsistently represented across different studies. Labeling each substrate individually would create a high-dimensional feature space, which is undesirable for a sparse dataset. To address this, we utilize OpenAI’s text-embedding models, specifically the text-embedding-3-small model, which converts input text into numerical vector representations. This embedding method generates a 128-dimensional vector, which is then reduced to 4- or 8-dimensional vectors using OpenAI’s API method, similar to principal component analysis. The text embedding is performed as follows 76:
@@
where $\mathbf{O}\in\mathbb{R}^{n\times v}$ is the input vector for $n$ tokens, $E\in\mathbb{R}^{v\times d}$ is the embedding matrix, and $\mathbf{V}\in\mathbb{R}^{n\times d}$ is the output embedding vector containing the embeddings for all tokens. Dimensions $v$, $n$, and $d$ correspond to the vocabulary size, number of tokens, and size of each token embedding vector, respectively. The embedding methodology implementation is elaborated in Supplementary Note 2 and Supplementary Figures 6 and 7. An alternative approach involves using ChatGPT to identify standard descriptors for substrate attribute based solely on the text information provided. This method can function independently but can also be combined with the embedding strategy to enhance feature creation. Table 3 summarizes the adopted feature engineering techniques, with further details in Supplementary Note 2. Supplementary Table 2 presents examples of standard descriptors generated using ChatGPT and converted into embedding vectors via text-embedding-3-small. The engineered features in the dataset are then discretized using the methods described in the following section.
@@ 2.5 Machine Learning Models
Here, we examine four different machine learning models to assess classification performance: Decision Tree, Random Forest, and XGBoost, which are all tree-based models, and Support Vector Machine (SVM), a kernel-based model. SVM has demonstrated strong learning capacity for classification tasks by separating classes through margin maximization, creating a decision boundary or hyperplane between positive and negative classes. Supplementary Note 4.2 provides details on the SVM formulation and optimization, while Supplementary Note 4 covers additional machine learning model details.
The collected data, once imputed, feature engineered and discretized, serve as training and test datasets for the classification task. In our dataset, about 30% and 19% of data points corresponding to bilayer and multi-layer graphene growth, respectively, while the remaining data ($\approx$ 50%) corresponds to monolayer graphene. We employ stratification when splitting the train-test data, which is a technique to ensure that the distribution of data across different classes of a subspace is representative of the original population (detailed in Supplementary Note 4.3). In this work, we use stratification primarily to create a representative set while calculating the learning curves using a 5-fold cross-validation.
@@ 2.6 Fine Tuning ChatGPT
ChatGPT-4o-mini was fine-tuned using OpenAI’s job creation tool, which allows the specification of training and validation datasets along with some hyperparameters. The goal of fine-tuning is to retrain a few layers of the LLM model such that it’s responses could be targeted to a desired output format and possibly to draw correlations from the input prompt patterns to predict the output. Specific details on the dataset preparation and fine-tuning models are given in Supplementary Note 5. We choose three epochs to calculate the average cross-validation score which is translated into token probability scores using a cross-entropy loss formula as described in Supplementary Note 5. In summary, the fine-tuning job involves (a) dataset preparation, (b) job creation, and (c) API call to the fine-tuned model. The API call retrieves classification results from our fine-tuned models, followed by accuracy report computation, similar to our approach with the SVM model. These fine-tuned models are compared to the SVM model by the accuracy, precision, recall, and F1 score.

### v2

@@ Abstract
Machine learning in materials science faces challenges due to limited experimental data, as generating synthesis data is costly and time-consuming, especially with in-house experiments. Mining data from existing literature offers a potential solution but introduces issues like mixed data quality, inconsistent formats, and variations in reporting experimental parameters, complicating the creation of consistent features for the learning algorithm. Additionally, combining continuous and discrete features can hinder the learning process with limited data. Here, we propose strategies that utilize large language models (LLMs) to enhance machine learning performance on a limited, heterogeneous dataset of graphene chemical vapor deposition synthesis compiled from existing literature. These strategies include prompting modalities for imputing missing data points and leveraging large language model embeddings to encode the complex nomenclature of substrates reported in chemical vapor deposition experiments. Imputed data from various prompting modalities are compared to statistical K-nearest neighbors imputation, with the former yielding a more diverse distribution, richer feature representation, and improved model generalization. We also demonstrate that discretizing continuous features enhances prediction accuracy. The proposed strategies enhance graphene layer classification using a support vector machine (SVM) model, increasing binary classification accuracy from 39% to 65% and ternary accuracy from 52% to 72%. We compare the performance of the SVM and a GPT-4 model, both trained and fine-tuned on the same data. Our results demonstrate that the numerical classifier, when combined with LLM-driven data enhancements, outperforms the standalone LLM predictor, highlighting that improving predictive learning with LLM strategies requires more than simple fine-tuning on datasets in data-scarce scenarios. Instead, it necessitates sophisticated approaches for data imputation and feature space homogenization to achieve optimal performance. The proposed strategies emphasize data enhancement techniques rather than merely refining learning architectures or regularizing loss functions, offering a broadly applicable framework for improving machine learning performance on scarce, inhomogeneous datasets.
@@ 1 Introduction
Large language models (LLMs) have recently emerged as a new approach in predictive chemistry and materials science 37, 38, 39, 40, 41, 42, 43, 44, 45. Specifically, they hold significant potential to address data scarcity and heterogeneity challenges, particularly by utilizing the extensive embedded pre-trained knowledge to impute missing data or homogenize inconsistencies in reported features. LLMs are referred to as near-universal generalists 46, positioning them as readily accessible alternatives to the aforementioned strategies in materials informatics. Furthermore, LLMs can be fine-tuned for specific problems with minimal prompt engineering, making them technically feasible for navigating complex feature spaces 47. This flexibility and generality offer potential solutions to both underfitting and overfitting challenges in machine learning on small datasets while also enabling LLMs to minimize hallucinations during data generation or featurization. In essence, LLMs could help address the bias-variance tradeoff that most ML strategies struggle to resolve when working with limited data. In this study, we demonstrate an LLM-assisted enhancement of machine learning using a sparsely populated and inherently heterogeneous dataset on chemical vapor deposition (CVD) of graphene, compiled from multiple studies in the literature.
@@
The CVD synthesis of graphene involves a wide range of experimental setups, with several key parameters governing the growth of graphene layers 48, 49, 50. To this date, no comprehensive dataset encompassing an exhaustive set of synthesis parameters has been developed for graphene CVD growth or synthesizing any 2D (atomic-layer-thick) material. Consequently, most machine learning studies on 2D material growth rely on limited, in-house experimental data, often constrained to specific ranges of synthesis conditions (e.g., pressures, temperatures, and precursor flow rates). These studies typically utilize machine learning models that require minimal feature engineering, as the in-house experimental data is already labeled and well-structured 51, 52, 53, 54, 55. For example, Hwang et al. developed a machine learning model to predict the morphology of graphene in the form of SEM (Scanning Electron Microscopy) images based on CVD processing variables (i.e., temperature, annealing time, growth time, and supply of hydrogen) 55. The size, coverage, domain density, and aspect ratio of CVD-grown graphene in SEM images were automatically measured using a region proposal convolutional neural network (R-CNN). A neural network and support vector machine were trained to map the relationship between CVD processing variables and these morphological features. Subsequently, a generative adversarial network (GAN) was designed to generate SEM images based on specified CVD processing conditions 55. However, this study relied on in-house experimental data, restricted to four processing variables and limited by fixed chamber conditions, including a constant precursor flow rate and the exclusive use of a single substrate type (Cu). Another notable effort by Schiller et al. involved creating a crowd-sourced database compiling an extensive list of parameters influencing various aspects of graphene CVD growth, contributed by multiple experimental groups 15. They also developed a UNet model for SEM image characterization, accessible as a nanoHub tool 56. However, the database remains incomplete, making it challenging to integrate it into a machine-learning framework for predicting outcomes based on CVD processing conditions.
@@
Recent advancements in large language models enable them to perform complex tasks such as understanding context and semantics and generating contextually relevant responses. The vast training data, pretraining on diverse corpora, and the sheer scale of these models make them excel at interpreting meaning and delivering accurate, coherent outputs. Several studies have utilized LLMs for imputation tasks 65, 66, 67, 68, 69, 70, 71, showing superior performance compared to traditional techniques such as K nearest neighbor (KNN), mean imputation, multivariate imputation by chained equations (MICE) and ML methods such as generative and discriminative methods (GAN) 72, 70, 66, 73. For example, Nazir et al. utilized ChatGPT-3.5 as an imputation engine to populate missing values for biological and psychological data, showcasing its out-performance compared to KNN and mean imputation65. They applied an iterative human-in-the-loop feedback process to refine ChatGPT’s imputation response by comparing the preceding response with ground-truth data, resulting in context-aware and accurate imputations of missing values in subsequent steps. Another study by Hayat et al.66 introduced an LLM-based imputation formalism to populate missing values for seven multivariate classification datasets from the UCI repository74. They generated missing-value descriptors using a pre-trained LLaMA2 model and then fine-tuned the model using the descriptor-enriched data set for classification tasks. The classification tasks were directly performed using the fine-tuned LLMs. Other notable studies have incorporated techniques such as QnA prompts for data imputation and property extraction using LLMs across a wide variety of tasks 73, 75, 72. This study focuses on leveraging the existing pre-trained ChatGPT-4o-mini for a general imputation strategy. Additionally, the learning capacities of numerical models, such as support vector machines and random forests, are evaluated for the target classification task and are compared with a fine-tuned ChatGPT-4o-mini counterpart.
@@
LLM-based data imputation is benchmarked against KNN-based imputation for the classification task. For the LLM-based approach, various prompt-engineering strategies guide the imputation task with specific instructions. After imputation, LLM-based featurization is used to unify and homogenize the numerical features for training binary and ternary classifiers, which predict the number of graphene layers based on a set of CVD processing inputs. We compare the effectiveness of different prompt-engineering strategies with traditional KNN-based data imputation, highlighting the superior performance of specific prompt-engineered GPT models. The prompt-driven imputed values are benchmarked with existing ground truth, and the missing values are directly imputed as an independent task. LLM-based strategies minimize imputation errors while enhancing data diversity, yielding a knowledge-driven, less imitative data distribution that outperforms KNN in machine learning. Additionally, we compare traditional non-LLM label-encoded features, commonly used for handling inconsistent labels, with LLM-generated features. Our results demonstrate the superior performance of LLM-embedded features in enhancing ML performance compared to label- or category-encoded features. Finally, we discretize the attributes to enhance the model’s learning performance compared to their continuous counterparts. These techniques improve ML performance for graphene layer classification, as detailed in the following sections.
@@ 2.1 Data Collection
Experimental literature was used to manually extract 164 data entries 57, 58, 59, 60, 50, 61, 62, 63, 64, each consisting of 10 attributes. Five of these attributes represent precursor flow rates (sccm), while the remaining chamber parameters include pressure (mbar), temperature (${}^{\circ}\text{C}$), growth time (min), and substrate. The target variable is the number of graphene layers. Many studies also report the CVD method; however, this primarily reflects the pressure range used in the CVD process. Therefore, we did not include it as a separate attribute. Table 1 shows a subset of the dataset used in this study. The dataset contains missing values for pressure, growth time, number of layers, and nearly all precursor flow rates. The precursor flow rates for each data entry are presented in the graphene_growth_conditions_layers.csv supplementary file. The raw data distributions across various attributes are shown in Supplementary Figure 1. More details on data collection are provided in Supplementary Note 1.
@@ 2.2 Imputation of Missing Attributes
Besides KNN imputation, we employed prompt-based LLM imputations to handle better sparsely populated or nearly empty attributes, such as the gas flow rates, which did not provide a well-defined data distribution for a distance-based algorithm like KNN. A more “generally-informed” model, such as an LLM, can leverage syntactical relationships between words to generate data that closely approximate values that might have been reported in the literature. Figure 2 shows the prompt-based LLM imputation methodology developed in this study. The prompts used for this task are named based on their general purpose as summarized in Table 2. Supplementary Note 1.1 details the methodology for each prompt. The GIST prompt introduced here is inspired by an approximation of the CLAIM prompt approach introduced in Ref.66. GIST differs from CLAIM as it only incorporates the language descriptors as a pre-processing step to assist the imputation and does not create a missingness-aware description of the data, which is central to the CLAIM methodology. Identifying missingness patterns is not feasible with high-dimensional, low-cardinality data like ours. The imputed datasets with each method are carried onto the next step of feature engineering. Supplementary Notes 1.3 to 1.5 provide details on imputation methods and their benchmarking.
@@
where LLM model = ChatGPT-4o-mini, context vector = [attributes] - [target], attribute = current attribute being imputed. We also demonstrate a benchmarking experiment where ChatGPT imitates the KNN imputation approach, shown in Supplementary Figure 3. Additional details about the nature of the imputed data points are shown in Supplementary Figures 4 and 5.
@@ 2.3 Feature Engineering
We conduct feature engineering to process the substrate attribute, the target variable, and the number of layers. The latter is handled using label encoding, classifying graphene layers into binary categories (monolayer versus multilayer) or ternary categories (monolayer versus bilayer versus multilayer), as the values are discrete. However, for the substrate attribute, as shown in Table 1, the data was inconsistently represented across different studies. Labeling each substrate individually would create a high-dimensional feature space, which is undesirable for a sparse dataset. To address this, we utilize OpenAI’s text-embedding models, specifically the text-embedding-3-small model, which converts input text into numerical vector representations. This embedding method generates a 128-dimensional vector, which is then reduced to 4- or 8-dimensional vectors using OpenAI’s API method, similar to principal component analysis. The text embedding is performed as follows 76:
@@
where $\mathbf{O}\in\mathbb{R}^{n\times v}$ is the input vector for $n$ tokens, $E\in\mathbb{R}^{v\times d}$ is the embedding matrix and $\mathbf{V}\in\mathbb{R}^{n\times d}$ is the output embedding vector containing the embeddings for all tokens. Dimensions $v$, $n$, and $d$ correspond to the vocabulary size, number of tokens, and size of each token embedding vector, respectively. The implementation of the embedding methodology is elaborated in Supplementary Note 2 and Supplementary Figures 6 and 7. An alternative approach involves using ChatGPT to identify standard descriptors for substrate attributes based solely on the text information provided. This method can function independently but can also be combined with the embedding strategy to enhance feature creation. Table 3 summarizes the adopted feature engineering techniques, with further details in Supplementary Note 2. Supplementary Table 2 presents examples of standard descriptors generated using ChatGPT and converted into embedding vectors via text-embedding-3-small. The engineered features in the dataset are then discretized using the methods described in the following section.
@@ 2.5 Machine Learning Models
Here, we examine four different machine learning models to assess classification performance: Decision Tree, Random Forest, and XGBoost, which are all tree-based models, and Support Vector Machine (SVM), a kernel-based model. SVM has demonstrated a strong learning capacity for classification tasks by separating classes through margin maximization and creating a decision boundary or hyperplane between positive and negative classes. Supplementary Note 4.2 provides details on the SVM formulation and optimization, while Supplementary Note 4 covers additional machine-learning model details.
The collected data, once imputed, feature-engineered, and discretized, serve as training and test datasets for the classification task. In our dataset, about 30% and 19% of data points correspond to bilayer and multi-layer graphene growth, respectively, while the remaining data ($\approx$ 50%) corresponds to monolayer graphene. We employ stratification when splitting the train-test data, which is a technique to ensure that the distribution of data across different classes of a subspace is representative of the original population (detailed in Supplementary Note 4.3). In this work, we use stratification primarily to create a representative set while calculating the learning curves using a 5-fold cross-validation.
@@ 2.6 Fine Tuning ChatGPT
ChatGPT-4o-mini was fine-tuned using OpenAI’s job creation tool, which allows the specification of training and validation datasets along with some hyperparameters. The goal of fine-tuning is to retrain a few layers of the LLM model such that its responses could be targeted to a desired output format and possibly to draw correlations from the input prompt patterns to predict the output. Specific details on the dataset preparation and fine-tuning models are given in Supplementary Note 5. We choose three epochs to calculate the average cross-validation score, which is translated into token probability scores using a cross-entropy loss formula as described in Supplementary Note 5. In summary, the fine-tuning job involves (a) dataset preparation, (b) job creation, and (c) API call to the fine-tuned model. The API call retrieves classification results from our fine-tuned models, followed by accuracy report computation, similar to our approach with the SVM model. These fine-tuned models are compared to the SVM model by the accuracy, precision, recall, and F1 score.